OpenStudy (anonymous):

What is the solution to the equation -3x - 4 = 5(x + 4)? someone plz help me :/

OpenStudy (***[isuru]***):

hi, -3x -4 = 5(x + 4) -3x -4 = 5x + 20 -4 - 20 =5x +3x 8x = -24 now u must be able to solve this one with ur eyes shut.... Wt will e the value of "x" then ?

OpenStudy (***[isuru]***):

did u get it ?

OpenStudy (anonymous):

no, idk what you mean?

OpenStudy (***[isuru]***):

did u get it up to the last step.. 8x = -24 ?

OpenStudy (anonymous):

yes

OpenStudy (***[isuru]***):

then (8x)/8 =(-24)/8 x = -3 did u get it now ?

OpenStudy (anonymous):

yes, thank you so much! you helped me a lot! :)

OpenStudy (***[isuru]***):

u r welcome!